The bank tells you that the intrest rate on a loan is 9% APR, if the compounding is monthly what is the effective rate? (state the answer in percentage terms 12.62%, 0.1262===> 12.62)
Answer #1

Effective rate = (1+(rate/period))^periods - 1

Effective rate = (1+(9%/12))^12 - 1

Effective rate = (1.0075)^12 - 1

Effective rate = 1.0938 - 1

Effective rate = 0.0938 or 9.38%